# Recent quantum runtime (dis)advantages

### v2

@@ II.3 Runtime for classical-quantum algorithms
Heuristic algorithms generally rely on a set of hyperparameters that must be tuned to achieve optimal performance [39]. Such tuning is often computationally expensive [40, 41]. If the reported runtime of a solver excludes the cost of hyperparameter optimization, it cannot be regarded as evidence of a runtime advantage. Unless the chosen hyperparameters are demonstrably problem-independent, or the tuning effort is uniform across all instances. Furthermore, in hybrid quantum–classical approaches, the reported runtime should also include the overhead associated with transferring data between the classical host and the quantum device. This practice is standard in classical–classical settings, such as combined CPU–GPU computations, and there is no justification for treating quantum–classical methods differently.
In this context, let us discuss the results of Ref. [42], where a new quantum-classical solver for large-scale spin glasses has been proposed, and it was claimed that it outperforms other quantum and classical algorithms both in terms of solution quality and the runtime. The solver is based on a sequential process, in which at every step the algorithm needs a set of hyperparameters. However, the runtimes presented by the authors do not explicitly include execution times needed to obtain those hyperparameters. Moreover, a tuning subroutine for hyperparameters is not described. Therefore, while the quality of solutions is impressive, the runtime advantage claim is not justified. Again, a proper metric in this context is $\mathrm{TT}\varepsilon$, with all the necessary time components included, and optimization performed [39].
@@ III Supremacy in the oracle query complexity framework
Here, we aim to interpret the query complexity framework from the perspective of a resource theory. To regard the number of oracle calls as a computational resource, one must assume that the oracle performs the most resource-intensive part of the computation. Only under this assumption the claim that an algorithm requiring fewer oracle queries is more efficient can be justified. Furthermore, even when a separation between the query complexities of different algorithms is proven, it is important to examine how this separation translates into runtime performance. This is particularly relevant in comparing quantum and classical computing, where the operations are characterized by different timescales – typically, a single quantum operation is slower than a classical one [48]. Because of these considerations, it has been argued that a practical quantum advantage can be achieved only by algorithms offering super-quadratic speedups [48]. Runtime analysis thus plays a crucial role in determining the problem scales at which an oracle query complexity advantage leads to a genuine computational advantage in practice.
Recently, it has been experimentally demonstrated that, for certain problem sizes and even in the presence of uncorrected noise, quantum computers can exhibit an exponential advantage in a variant of Simon’s problem [20]. In the original problem, the task is to find a hidden period (a bitstring) encoded in an unknown 22-to-11 function. In contrast, the authors of [20] study $N$-bit periods with a restricted Hamming weight $w$ (i.e., the number of nonzero bits). In their formulation, the algorithm’s figure of merit is defined in terms of a score function, whose primary purpose is to penalize random period-guessing strategies. The authors study scaling of this figure of merit, as a function of the total number of possible periods denoted as $N_{w}\equiv\sum_{j}\binom{N}{j}$, where $N$ is the total number of bits, and $w$ the maximal allowed Hamming weight. Theoretical considerations are presented that suggest possibility of a quantum advantage for such defined problem even for noise, uncorrected quantum computers. The exponential advantage manifest itself in a polylogarithmic scaling of the score function as a function of the total number of possible periods (the classical model scales polynomial with this parameter). The authors verify then these claims experimentally. An exponential speedup is observed for functions with $29$-bit inputs and restricted Hamming weights in the range $w\in[2,7]$. For larger problem sizes, however, the advantage is destroyed by noise. This raises the question: how does this favorable scaling relate to the actual runtime of the algorithm?
@@
FIG. 2. : For the restricted Simon’s problem the quantum advantage manifests itself in a favorable polylogarithmic scaling of the protocol score function with the total number of periods $N_{w}=\sum_{j}\binom{N}{j}$ as compared to the exponential scaling for the classical algorithm. The score function depends on number of oracle queries and a probability of protocol success. (a) Total number of oracle queries for the protocol and (b) runtime as a function of total number of bits $N$, as obtained via execution of classical and quantum algorithm solving restricted Simon’s problem with $w=7$. The shaded area indicates sizes of problems, for which the advantage was verified experimentally in [20]. For the considered problem sizes and oracle periods the classical algorithm runtime is shorter than the quantum one, we predict that the runtime crossover occurs for problem sizes $N=60$. For example, in the case of $N=29$ bits, what corresponds to the implementation presented in Ref. [20], the runtime of the classical algorithm is two order of magnitude shorter than the quantum algorithm. The quantum implementation is based on oracle constructed in Ref. [20], and the number of shots required for the algorithm was established using Eq. (16) of Ref. [20]. In this case the quantum circuit was transpiled to take into account connectivity and native gate set of IBM Brisbane, and runtime was estimated using Qiskit functionalities. The classical algorithm was implemented on a GPU. The fitting parameters of the score function for quantum and classical algorithm are presented in panels (c), (d). Parameter $\gamma=0$ indicates that the score scales polylogarithmicly, which implies quantum scaling advantage. The data for the quantum algorithm correspond to the results obtained for IBM Brisbane with dynamical decoupling - Table $X$ in [20].
@@ IV Appropriate choice of a classical reference algorithm
For this figure of merit, an optimal reference classical algorithm should ideally combine solution quality with short runtime. Thus, the most reasonable choices are parallelizable algorithms, whose execution times are significantly shorter than those of sequential approaches, thanks to their ability to fully exploit modern computing resources such as GPUs [53] or FPGAs [54]. In this regard, a particularly interesting class of classical heuristics is based on the dynamics of Hamiltonian nonlinear systems [35, 37, 36, 55, 56, 25, 57], as these methods combine good performance with short execution times. Recently, it was shown that one algorithm from this class-the Simulated Bifurcation Machine (SBM)-achieves scaling of $\mathrm{TT}\varepsilon$ comparable to, or better than, that of quantum annealing [25]. This result effectively closes the previously reported quantum-classical gap in approximate optimization announced in [19], where the chosen classical reference was Parallel Tempering with Isoenergetic Cluster Moves (PT-ICM). The PT-ICM algorithm belongs to the class of physics-inspired methods based on temperature annealing [58]. However, unlike algorithms rooted in nonlinear Hamiltonian dynamics, PT-ICM is not straightforwardly parallelizable, and consequently exhibits longer execution times. This makes it a suboptimal reference for comparison with quantum annealing, a conclusion supported by the detailed analysis in [25].
Secondly, a classical reference algorithm should be designed to solve the same class of problems. Otherwise, the comparison is biased as it favours the quantum algorithm. Potential worst performance of a chosen classical algorithm is not an evidence of a quantum algorithm advantage but it merely reflects the fact that the problem cannot be solved natively by the classical reference. For example, in [21], the investigated counter-diabatic quantum optimization algorithm is designed to solve HUBO problems, whereas the reference CPLEX solver requires reformulation into a mixed-integer program. As the authors note, this reformulation is responsible for the longer runtime of CPLEX compared to their quantum algorithm. Setting aside the fact that the overheads of counter-diabatic quantum optimization were not accounted for (see the discussion in Sec. II), the relevant question is how the performance of the counter-diabatic quantum optimization algorithm would compare against a classical solver capable of solving HUBO with significantly lower overhead?
@@
where $G_{2}$ and $G_{3}$ are the sets of two- and three-body couplings, respectively, and $J_{mn}$ and $K_{pqr}$ are the corresponding coupling strengths, The topology of considered instances is constructed iteratively, starting from the coupling graph $C_{0}$ of heavy-hexagonal lattice of IBM’s Heron architecture. Each step starts by using graph coloring to identify sets of independent two- and three-body interactions, then $S_{2q}$ ($S_{3q}$) of them gets included in the $G_{2}$ ($G_{3}$) set, and finally the coupling graph is modified by performing SWAP operation on pairs of qubits as defined by one of the two-body interactions sets. This procedure creates rather challenging instances, which are however particularly well suited for the BF-DCQO algorithm, since by changing the number of SWAP iterations, one can directly control the depth of the quantum circuit necessary for the quantum part of the algorithm. The authors of Ref. [21] thus restricted their analysis to only one SWAP iteration, which results in shallow circuits that can be executed on an IBM device in a regime, where the results are not dominated by noise. The instances used in Ref. [21] were not made publicly available  [^1] The authors did not respond to our request to share their HUBO instances. Consequently, we generated our own instances, closely following the description in Ref. [21] (cf. App. A3), and we make both these instances and the generation routine available in the accompanying GitHub [60].
@@
This analysis illustrates how important it is to select a suitable classical reference algorithm when assessing quantum advantage claims. Let us also mention that the results of [21] were recently also disputed in [61], where it was claimed that the superior performance of BF-DCQO is not due to the quantum routine, as a modified algorithm with the quantum routine replaced by a classical solver exhibits a similar performance. Finally, for the sake of completeness, we tested the performance of D-Wave’s newest quantum annealer, the Advantage2 1.6 based on the Zephyr topology. The results are presented in Fig. IV. Since the considered instances are not native to the topology of the QPU, an embedding procedure must be carried out. Fig. IV(c) shows the distribution of chain lengths in the embedding, with a tail extending to very long chains. This, unsurprisingly, results in a very poor performance, even after cherry-picking the best result out of $35$ random instances, and $5$ independent shots of $2^{10}$ samples each. We stress that this is just a single manifestation of a more general issue regarding claims of quantum advantage — the necessity of embedding problem instances onto the hardware graph of a quantum device, often significantly reduces the performance, and thus one should be very careful when making such claims on the basis of experiments with hardware-native problems only.
@@ VI Summary
Quantum computing has the potential to become one of the next disruptive technologies. However, realizing this potential requires not only technological progress but also the identification of (potentially industry-relevant) use cases where quantum computing can deliver a significant advantage over existing classical methods. In this work, we argue that such an advantage should be reflected in a shorter time-to-solution achieved by a quantum algorithm. From this perspective, we (re)examined recent claims of quantum advantage and found that, under a proper definition of runtime and with an appropriate choice of the classical reference, none of them demonstrate a genuine runtime supremacy. In particular, we analyzed two recent milestone results:
Quantum Annealing (approximate QUBO). Prior work [19] proxied runtime by the annealing time. However, a careful runtime measurements show the median $\mathrm{TT}\varepsilon$ is essentially flat over the tested sizes. This is due to the readout per shot being $\approx 200\mu s$ while the annealing durations are only $0.5\text{–}27\mu s$. Measurement errors are too large to draw meaningful conclusions; thus no scaling can be inferred, cf. Fig. II.1. Furthermore, this “supremacy” would not survive against stronger classical baseline (such as SBM) even if the annealing time would faithfully reflect the device’s runtime, as shown in [25]. SBM achieves robust scaling far beyond the system sizes attainable by current quantum annealers, cf. Fig. IV.
@@
Gate-based oracle query (restricted Simon’s problem). Although there is an exponential separation in query-complexity scaling, wall-clock performance favors a tuned classical GPU baseline: for $N=29$, and the “restriction” $w=7$ [20], classical runtime is $\approx 0.05\mathrm{s}$, and the quantum $\approx 2\mathrm{s}$ ($\sim 100\times$ slower). The projected crossover of classical and quantum runtimes is at $N\approx 60$ lies beyond the current capabilities of noise-limited devices, cf. Fig. III.
BF-DCQO hybrid algorithm. Finally, we scrutinized reported gains for solving HUBO [21] and show they are based on estimates that neglect robust timing protocols and cherry-picking statistics. With such methodology one can produce arbitrary supremacy by design, cf. Fig. IV. It is worth noting that the problem instances considered in [21] were specifically designed to align with IBM’s Heron architecture [62], and purportedly to be “challenging” for classical methods. We found that on those instances, the high-quality physics-inspired methods produces diverse solutions and therefore requires much richer statistics to substantiate any claim of supremacy. Cherry-picking the best out of a few runs ($5$ therein) is not sufficient, cf. App. A3.
@@
In conclusion, across all cases, a practical runtime advantage on current NISQ hardware remains elusive. Credible claims should not omit dominant overheads (readout, thermalization, transpilation, etc.), must employ a proper metric such as $\mathrm{TT}\varepsilon$ with all the required optimizations, and should benchmark against state-of-the-art, parallel classical solvers impartially selected on a case-by-case basis reflecting the problem class.
In contrast, in [19], all advantages were given to quantum annealing (such as avoiding dense instances and thus eliminating the need for embedding, allowing a relatively large $\epsilon>1\%$, and using annealing time as a proxy for the actual runtime), yet this approach still could not compete for at least “limited supremacy” with a highly optimized classical algorithm executed on a single GPU [25].

### v3

@@ II.3 Runtime for classical-quantum algorithms
Heuristic algorithms typically depend on a set of hyperparameters that must be tuned in order to achieve optimal performance [40]. Such tuning can be computationally demanding [41, 42]. As a result, reported solver runtimes that exclude the cost of hyperparameter optimization are difficult to interpret as evidence of a runtime advantage, unless the chosen hyperparameters can be shown to be problem-independent or the tuning effort is applied uniformly across all instances. In addition, for hybrid quantum–classical approaches, runtime accounting should include the overhead associated with data transfer between the classical host and the quantum device. This practice is standard in classical heterogeneous computing, such as CPU–GPU workflows, and analogous considerations apply in the quantum–classical context.
Within this framework, we consider the results of Ref. [43], where a quantum–classical solver for large-scale spin-glass problems was proposed and reported to achieve improved solution quality and favorable runtime performance relative to other approaches. The solver operates as a sequential procedure in which each step requires a set of hyperparameters. However, the reported runtimes do not explicitly include the time required to determine these hyperparameters, and the procedure used for their optimization is not described. Consequently, while the reported solution quality is notable, the available data do not allow for a fully reproducible assessment of end-to-end runtime performance.
From a benchmarking perspective, a more informative comparison would again rely on the time-to-epsilon metric $\mathrm{TT}\varepsilon$, with all relevant time contributions— including hyperparameter optimization—consistently accounted for, as advocated in Ref. [40].
@@ III Supremacy in the oracle query complexity framework
Here, we aim to interpret the query complexity framework from the perspective of a resource theory. To regard the number of oracle calls as a computational resource, one must assume that the oracle performs the most resource-intensive part of the computation. Only under this assumption can the claim that an algorithm requiring fewer oracle queries is more efficient be justified. Furthermore, even when a separation between the query complexities of different algorithms is proven, it is important to examine how this separation translates into runtime performance. This is particularly relevant in comparing quantum and classical computing, where the operations are characterized by different timescales – typically, a single quantum operation is slower than a classical one [49]. Because of these considerations, it has been argued that a practical quantum advantage can be achieved only by algorithms offering super-quadratic speedups [49]. Runtime analysis thus plays a crucial role in determining the problem scales at which an oracle query complexity advantage leads to a genuine computational advantage in practice.
Recently, it has been experimentally demonstrated that, for certain problem sizes and even in the presence of uncorrected noise, quantum computers can exhibit an exponential advantage in a variant of Simon’s problem [21]. In the original problem, the task is to find a hidden period (a bitstring) encoded in an unknown 22-to-11 function. In contrast, the authors of [21] study $N$-bit periods with a restricted Hamming weight $w$ (i.e., the number of nonzero bits). In their formulation, the algorithm’s figure of merit is defined in terms of a score function, whose primary purpose is to penalize random period-guessing strategies. The authors study scaling of this figure of merit, as a function of the total number of possible periods denoted as $N_{w}\equiv\sum_{j}\binom{N}{j}$, where $N$ is the total number of bits, and $w$ the maximal allowed Hamming weight. Theoretical considerations are presented that suggest the possibility of a quantum advantage for such defined problem even for noisy, uncorrected quantum computers. The exponential advantage manifest itself in a polylogarithmic scaling of the score function as a function of the total number of possible periods (the classical model scales polynomially with this parameter). The authors then verify these claims experimentally. An exponential speedup is observed for functions with $29$-bit inputs and restricted Hamming weights in the range $w\in[2,7]$. For larger problem sizes, however, the advantage is destroyed by noise. This raises the question: how does this favorable scaling relate to the actual runtime of the algorithm?
@@
FIG. 2. : For the restricted Simon’s problem the quantum advantage manifests itself in a favorable polylogarithmic scaling of the protocol score function with the total number of periods $N_{w}=\sum_{j}\binom{N}{j}$ as compared to the exponential scaling for the classical algorithm. The score function depends on number of oracle queries and a probability of protocol success. (a) Total number of oracle queries for the protocol and (b) runtime as a function of total number of bits $N$, as obtained via execution of classical and quantum algorithm solving restricted Simon’s problem with $w=7$. The shaded area indicates sizes of problems, for which the advantage was verified experimentally in [21]. For the considered problem sizes and oracle periods the classical algorithm runtime is shorter than the quantum one, we predict that the runtime crossover occurs for problem sizes $N=60$. For example, in the case of $N=29$ bits, what corresponds to the implementation presented in Ref. [21], the runtime of the classical algorithm is two orders of magnitude shorter than the quantum algorithm. The quantum implementation is based on oracle constructed in Ref. [21], and the number of shots required for the algorithm was established using Eq. (16) of Ref. [21]. In this case the quantum circuit was transpiled to take into account connectivity and native gate set of IBM Brisbane, and runtime was estimated using Qiskit functionalities. The classical algorithm was implemented on a GPU. The fitting parameters of the score function for quantum and classical algorithm are presented in panels (c), (d). Parameter $\gamma=0$ indicates that the score scales polylogarithmically, which implies quantum scaling advantage. The data for the quantum algorithm correspond to the results obtained for IBM Brisbane with dynamical decoupling - Table $X$ in [21].
@@ IV Appropriate choice of a classical reference algorithm
For this figure of merit, an optimal reference classical algorithm should ideally combine solution quality with short runtime. Thus, the most reasonable choices are parallelizable algorithms, whose execution times are significantly shorter than those of sequential approaches, thanks to their ability to fully exploit modern computing resources such as GPUs [54] or FPGAs [55]. In this regard, a particularly interesting class of classical heuristics is based on the dynamics of Hamiltonian nonlinear systems [36, 39, 37, 56, 57, 26, 58], as these methods combine good performance with short execution times. Recently, it was shown that one algorithm from this class—the Simulated Bifurcation Machine (SBM)-achieves scaling of $\mathrm{TT}\varepsilon$ comparable to, or better than, that of quantum annealing [26]. This result effectively closes the previously reported quantum-classical gap in approximate optimization announced in [20], where the chosen classical reference was Parallel Tempering with Isoenergetic Cluster Moves (PT-ICM). The PT-ICM algorithm belongs to the class of physics-inspired methods based on temperature annealing [59]. However, unlike algorithms rooted in nonlinear Hamiltonian dynamics, PT-ICM is not straightforwardly parallelizable, and consequently exhibits longer execution times. This makes it a suboptimal reference for comparison with quantum annealing, a conclusion supported by the detailed analysis in [26]. Note that a detailed discussion of the performance of a class of meta-heuristic classical algorithms for approximate optimization on cubic-lattice tile-planting models is presented in [60].
Secondly, a classical reference algorithm should be designed to solve the same class of problems. Otherwise, the comparison is biased as it favours the quantum algorithm. The potentially worse performance of a chosen classical algorithm is not evidence of a quantum algorithm advantage but it merely reflects the fact that the problem cannot be solved natively by the classical reference. For example, in [22], the investigated counter-diabatic quantum optimization algorithm is designed to solve HUBO problems, whereas the reference CPLEX solver requires reformulation into a mixed-integer program. As the authors note, this reformulation is responsible for the longer runtime of CPLEX compared to their quantum algorithm. Setting aside the fact that the overheads of counter-diabatic quantum optimization were not accounted for (see the discussion in Sec. II), the relevant question is how the performance of the counter-diabatic quantum optimization algorithm would compare against a classical solver capable of solving HUBO with significantly lower overhead?
@@
where $G_{2}$ and $G_{3}$ are the sets of two- and three-body couplings, respectively, and $J_{mn}$ and $K_{pqr}$ are the corresponding coupling strengths. The topology of considered instances is constructed iteratively, starting from the coupling graph $C_{0}$ of heavy-hexagonal lattice of IBM’s Heron architecture. Each step starts by using graph coloring to identify sets of independent two- and three-body interactions, then $S_{2q}$ ($S_{3q}$) of them gets included in the $G_{2}$ ($G_{3}$) set, and finally the coupling graph is modified by performing SWAP operation on pairs of qubits as defined by one of the two-body interactions sets. This procedure creates rather challenging instances, which are however particularly well suited for the BF-DCQO algorithm, since by changing the number of SWAP iterations, one can directly control the depth of the quantum circuit necessary for the quantum part of the algorithm. The authors of Ref. [22] thus restricted their analysis to only one SWAP iteration, which results in shallow circuits that can be executed on an IBM device in a regime, where the results are not dominated by noise. The instances used in Ref. [22] were not made publicly available  [^1] The authors did not respond to our request to share their HUBO instances. Consequently, we generated our own instances, closely following the description in Ref. [22] (cf. App. A3), and we make both these instances and the generation routine available in the accompanying GitHub [62].
@@
This analysis illustrates how important it is to select a suitable classical reference algorithm when assessing quantum advantage claims. Let us also mention that the results of [22] were recently also disputed in [63], where it was claimed that the superior performance of BF-DCQO is not due to the quantum routine, as a modified algorithm with the quantum routine replaced by a classical solver exhibits a similar performance. Finally, for the sake of completeness, we tested the performance of D-Wave’s newest quantum annealer, the Advantage2 1.6 based on the Zephyr topology. The results are presented in Fig. IV. Since the considered instances are not native to the topology of the QPU, an embedding procedure must be carried out. Fig. IV(c) shows the distribution of chain lengths in the embedding, with a tail extending to very long chains. This, unsurprisingly, results in a very poor performance, even after post-selecting the best result out of $35$ random instances, and $5$ independent shots of $2^{10}$ samples each. We stress that this is just a single manifestation of a more general issue regarding claims of quantum advantage — the necessity of embedding problem instances onto the hardware graph of a quantum device, often significantly reduces the performance, and thus one should be very careful when making such claims on the basis of experiments with hardware-native problems only.
@@ VI Summary
Quantum computing has the potential to become a disruptive technology. Realizing this potential, however, requires not only continued technological progress but also the identification of use cases – potentially relevant for industry – in which quantum methods provide a clear advantage over existing classical approaches. In this work, we adopt a pragmatic perspective in which such an advantage should manifest as a shorter end-to-end time-to-solution for a quantum algorithm.
From this perspective, we re-examined several recent claims of quantum advantage and find that, when runtime is defined operationally and compared against appropriately chosen classical references, none of the considered demonstrations establish a practical runtime advantage on current NISQ hardware. We focus in particular on three representative case studies.
Quantum annealing (approximate QUBO). In Ref. [20], runtime was proxied by the annealing time. Using end-to-end runtime measurements, we find that the median $\mathrm{TT}\varepsilon$ remains approximately constant over the tested problem sizes. This behavior is dominated by the readout time per shot ($\sim 200\,\mu\mathrm{s}$), which exceeds the annealing duration ($0.5\text{--}27\,\mu\mathrm{s}$) by one to two orders of magnitude. As a result, statistical uncertainties are too large to support reliable scaling conclusions over the explored regime (cf. Fig. II.2). Moreover, comparison with a stronger classical baseline, such as the Simulated Bifurcation Machine (SBM), indicates that no runtime advantage is observed even under optimistic assumptions regarding the annealing time, consistent with Ref. [26]. SBM exhibits robust scaling behavior at system sizes well beyond those accessible to current quantum annealers (cf. Fig. IV).
@@
Gate-based oracle query (restricted Simon’s problem). While an exponential separation in query complexity is present, wall-clock performance in the experimentally accessible regime favors a tuned classical GPU implementation. For $N=29$ and restriction parameter $w=7$ [21], the classical runtime is approximately $0.05\,\mathrm{s}$, whereas the quantum implementation requires roughly $2\,\mathrm{s}$, corresponding to a $\sim 100\times$ difference. Extrapolations suggest a crossover only at $N\approx 60$, which lies beyond the capabilities of current noise-limited devices (cf. Fig. III).
BF-DCQO hybrid algorithm. Finally, we analyze reported performance gains for solving HUBO instances [22]. We find that these conclusions rely on runtime estimates and limited statistical sampling, which complicate a reproducible assessment of end-to-end performance. In particular, the problem instances considered were designed to align with specific hardware characteristics [64] and require substantially richer statistics to support claims of systematic advantage. Our analysis shows that, under consistent timing protocols and broader sampling, the reported performance differences are not robust (cf. Fig. IV and App. A3).
@@
In summary, across all examined cases, a practical runtime advantage on current NISQ hardware is not observed under experimentally grounded performance metrics. Credible claims of runtime quantum advantage require careful accounting of dominant overheads (including readout, thermalization, and compilation), the use of well-defined metrics such as $\mathrm{TT}\varepsilon$ with appropriate optimization, and benchmarking against state-of-the-art classical solvers selected impartially for the problem class under consideration.
VII Outlook
No single, universal procedure exists for evaluating quantum advantage, since different problem classes and computational paradigms exhibit distinct characteristics. Nevertheless, a set of general principles can be distilled from our analysis. In the context of runtime-based quantum advantage, we summarize these principles as follows:
  • Primary Runtime Measurement: Runtime measurements should be obtained from direct timing of executions, rather than inferred from nominal device parameters or documentation.
  • Comprehensive Overhead Accounting: Runtime should be reported as a direct measurement of the total execution time, including all relevant system-level overheads such as readout, compilation, data transfer, and hyperparameter tuning.
  • Standardized Performance Metrics: Performance should be evaluated using well-defined and widely accepted metrics, such as the time-to-$\varepsilon$ for sufficiently small values of $\varepsilon$, with results assessed across diverse problem instances and structures.
  • Scaling Integrity: Scaling analyses should span a sufficiently broad range of problem sizes to capture asymptotic behavior, and extrapolations beyond the measured regime should be treated with appropriate caution.
  • High-Performance Classical Baselines: Classical reference algorithms should be selected with care, taking into account the current state of algorithmic development and implementation-level optimizations.
  • The Hybrid Null Hypothesis: For hybrid quantum–classical algorithms, a suitable null hypothesis should be tested, namely whether replacing the quantum subroutine with a purely classical counterpart yields comparable or improved performance.
We stress that these principles are intended as general guidance rather than problem-specific prescriptions, and are not meant to constitute an exhaustive set of criteria.
We emphasize that while energy efficiency is a potentially relevant dimension of quantum advantage [14], the present work deliberately focuses on runtime-based performance metrics. At the current level of user access to quantum hardware, energy consumption cannot be independently verified or consistently defined, as it depends on system-level factors such as cryogenics and control electronics that are typically not transparent to the user. At the same time, contemporary high-performance classical accelerators, such as modern GPUs, typically operate at sub-kilowatt power levels, underscoring the importance of clearly defined system boundaries when comparing different computational platforms. The development of verifiable and reproducible energy-efficiency metrics therefore remains an important open challenge for future benchmarking efforts.
As a final note, while the present work does not focus on system-level capability metrics such as Quantum Volume [65] and CLOPS [66], we note that such metrics provide a valuable and complementary means of tracking technological progress in quantum hardware, particularly at the level of device capability and system throughput.